# LLM-Card: Towards a Description and Landscape of Large Language Models

### v2

@@ Abstract
With the rapid growth of the Natural Language Processing (NLP) field, a vast variety of Large Language Models (LLMs) continue to emerge for diverse NLP tasks. As more papers are presented, researchers and developers face the challenge of information overload. Thus, it is particularly important to develop a system that can automatically extract and organise key information about LLMs from academic papers (LLM model card). This work is to develop such a pioneer system by using Named Entity Recognition (NER) and Relation Extraction (RE) methods that automatically extract key information about large language models from the papers, helping researchers to access information about LLMs efficiently. These features include model licence, model name, and model application. With these features, we can form a model card for each paper. Data-contribution wise, 106 academic papers were processed by defining three dictionaries - LLMs name, licence, and application. 11,051 sentences were extracted through dictionary lookup, and the dataset was constructed through manual review of the final selection of 129 sentences with a link between the name and the licence, and 106 sentences with a link between the model name and the application. Data and code in LLM-Card is openly hosted at https://github.com/shengwei-tian/dependency-parser-visualization
1 Introduction
@@
As a major approach in natural language processing (NLP), language modelling has been widely studied for language understanding and generation in the past two decades, evolving from statistical language models to neural language models. Recently, Pre-trained Language Models (PLMs) have been proposed by pretraining Transformer models over large-scale corpora, showing strong capabilities in solving various NLP tasks. Since researchers have found that model scaling can lead to an improved model capacity, they further investigate the scaling effect by increasing the parameter scale to an even larger size. Interestingly, when the parameter scale exceeds a certain level, these enlarged language models not only achieve a significant performance improvement but also exhibit some special abilities (e.g., in context learning) that are not present in small-scale language models (e.g., BERT). To discriminate the language models in different parameter scales, the research community has coined the term LLMs for the PLMs of significant size e.g., tens or even hundreds of billions of parameters.
In the vast world of computer language processing, a revolutionary entity has emerged, Large Language Models (LLMs). These models possess the remarkable ability to understand intricate linguistic patterns and respond coherently matching the context. LLMs are a type of Artificial Intelligence (AI) that serve as powerful tools for various tasks, including machine translation and question answering. One notable example of LLMs is ChatGPT, developed by OpenAI. Based on the GPT-3.5 architecture, it was trained on a vast corpus of internet-sourced text data, including books, articles, wikis, and websites. ChatGPT excels at generating human-like responses and engaging in conversations with users Hadi et al. (2023).
@@
In recent years, major research institutes and technology companies have increased their investment in the development of LLMs, resulting in a series of groundbreaking models with the potential for a wide range of applications. OpenAI is a pioneer in this field. The GPT family (e.g., GPT-3 and GPT-4) has made significant advances in natural language generation, dialogue systems, and text comprehension, and is widely used in a variety of intelligent assistants, content creation tools, and automated services Torfi et al. (2020). Meanwhile, the Bidirectional Encoder Representations from Transformers (BERT) Devlin et al. (2019a) and the subsequent T5 (Text-to-Text Transfer Transformer) Raffel et al. (2020) models developed by Google research department have also made significant progress in tasks such as question-answer systems, text classification and information retrieval. Information Retrieval and BERT is used in the natural language processing of Google’s search engine. The Roberta and BlenderBot series of models from Facebook (now Meta) have demonstrated powerful performance in social media data analysis and conversational AI Liu (2019).
1.1 Motivation
@@
Natural Language Processing (NLP) technologies, driven by Large Language Models (LLMs), have skyrocketed and revolutionised the field of NLP, leading to breakthroughs in various tasks such as machine translation, text summarisation, and dialogue systems. However, these advances have been accompanied by the challenge of navigating the vast array of recent research findings. Each year, thousands of papers on LLMs are published, presenting new models, methods, and innovations. This leads to a phenomenon of information overload that makes it difficult for even the most experienced researchers to keep up with the latest developments.
Figure 1: LLM-card method demonstration
@@
We need a convenient and efficient tool that enables researchers and developers to quickly sift through and understand large numbers of models and their properties, i.e. the LLM-card. Motivated by these needs, this work proposes a novel system that automatically extracts such relationships, as illustrated in Figure 1. Our goal is to simplify the process of understanding key information about licence types, application areas, and more for researchers. By automating this process, researchers can gain a clearer understanding of the latest models, allowing them to save time to focus on innovation rather than tediously sifting through extensive literature.
1.2 Research Questions
@@
1.3 Highlights
To answer these Research Questions (RQs), we have the following deliverables and highlights regarding this investigation.
  • Define a dictionary look-up method to identify the targeted sentences via key phrases.
  • Explore a rule-based method to map the LLMs and their model cards, such as licences and applications.
  • Manually construct annotated data on such a task and contribute this as the first public data set in this field.
1.4 Paper Structure
The remainder of this paper is structured as follows:
  • Section 2: Literature Review - This section provides an in-depth review of the current research and developments in the field of LLMs, highlighting existing key models, architectures, and methodologies. It also discusses the challenges associated with managing the vast amount of information in this rapidly evolving field.
  • Section 3: Methodology - This section outlines methodologies employed in developing the automated information extraction system, including data collection, processing techniques, and system design.
  • Section 4: Experiment - This section presents the results of our system implementation, including the comparison tables generated by the system, and discusses the implications of these results in the context of current research and industry needs.
  • Chapter  5: Conclusions and Future Work - This section summarises the findings of our investigation, discusses the limitations of the current system, and proposes directions for future research and development.
2 Literature Review
2.1 Large Language Models
Language has always been one of the most important abilities that human beings rely on for survival, and this ability develops in childhood and evolves throughout life. Machines, however, lack this innate language ability. So, how can we enable machines this ability to understand and process human language, so that machines can read, communicate, and write like human beings? This is a long-term challenge that humanity continues to face Turing (2009).
Early NLP approaches largely relied on rule-based systems and statistical methods, which, despite their effectiveness, were limited by their dependence on handcrafted features and domain-specific knowledge. The introduction of neural networks in the 1980s began to shift this landscape, but it was the advent of deep learning and the subsequent development of LLMs that truly revolutionised the field.
A pivotal moment in the history of LLMs came with the development of word embeddings, such as word2vec Mikolov (2013), which enabled models to learn distributed representations of words in a continuous vector space. This innovation laid the groundwork for more complex architectures like Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks, which were among the first to effectively model sequential data, capturing dependencies across different points in a text sequence.
However, the real breakthrough in LLMs occurred with the introduction of the Transformer architecture Vaswani (2017). Unlike RNN, which processed sequences sequentially, Transformers could process all tokens in parallel, thanks to their self-attention mechanism. This innovation dramatically improved the efficiency and scalability of language models, enabling them to handle larger datasets and more complex tasks. The first major success using Transformers was Google’s BERT (Bidirectional Encoder Representations from Transformers), which was introduced in 2018. BERT was revolutionary in its bidirectional approach to text encoding, allowing the model to consider the context from both sides of a word, rather than just the preceding text as in traditional models Devlin et al. (2019b).
Following BERT, the field of NLP has progressed rapidly with a series of more powerful models, notably OpenAI’s GPT family. GPT-2, released in 2019, significantly improves AI text generation and closes the gap with human writing. In 2020, GPT-3 set a new benchmark with 175 billion parameters for performing small or no samples for complex tasks such as code writing and dialogue generation Brown (2020). In 2022, ChatGPT based on GPT-3.5 further optimises dialogue performance for more natural interactions. GPT-4 in 2023 extends to multi-modal tasks, handling text and image inputs, showing potential for a wide range of applications in fields such as healthcare and law OpenAI (2023).
@@ 2.1 Large Language Models
2.2 Named Entity Recognition
@@
Named Entity Recognition (NER) is a subtask of information extraction in NLP, which is used to classify named entities into predefined categories such as names of people, organisations, places, etc. In the field of NLP, understanding these entities is crucial for many applications because they often contain the most important information in a text. NER is a bridge between unstructured text and structured data, enabling machines to sift through large amounts of textual information and extract valuable data in the form of classification. By precisely locating specific entities in a sea of words, NER changes the way we process and utilise textual data. In text mining, a named entity is a word or a phrase that clearly identifies an entity, an entity with similar properties, from a set of other items with similar properties. In the formulation of named entities, the word named limits the scope of entities that have one or more rigid designators that represent the referent. Typically, rigid designators often include proprietary names, but this depends on the domain of interest, which has reference words for objects as named entities Sharnagat (2014). For example, in the sentence: BERT was released under the Apache licence 2.0, the entity of licence is Apache licence and the entity of the model name is BERT.
Figure 2: The steps for named entities recognition
@@
As shown in Figure 2, NER can be broken down into several steps: Tokenisation before identifying entities, the text is split into tokens, which can be words, phrases, or even sentences. For example: BERT was released under the Apache licence 2.0 would be split into word chunks such as ‘BERT’, ‘was’, ‘released’, ‘under’, ‘the’, ‘Apache’, ‘licence’, ‘2.0’. Entity identification involves recognising items such as capitalisation in licence (‘Apache 2.0’) or specific formats. Once entities are identified, they are grouped into predefined categories such as ‘model name’, ‘licence’ or ‘application’. This is typically achieved by machine learning models trained on annotated datasets. In our example, ‘BERT’ would be categorised as ‘model name’, while ‘Apache 2.0’ would be categorised as ‘licence’. The NER system usually considers the surrounding context to improve accuracy. After the initial identification and classification, post-processing may be performed to refine the results, which may involve resolving ambiguities, merging multi-tagged entities, or augmenting the entity data with a knowledge base Jurafsky (2000).
2.3 Relation Extraction
2.3.1 Rule-based methods
Figure 3: Rule-based methods for Relation Extractions
@@ 2.3.1 Rule-based methods
The rule-based relation extraction approach as shown in Figure 3 is one of the early techniques in the field of NLP, which relies on linguists and domain experts to manually write rules to extract relations from text based on an in-depth understanding of linguistic structures Zhang et al. (2009). The core of this approach lies in the use of predefined linguistic patterns or syntactic structures to accurately match entities and their relationships in a text. While this approach excels in some specific domains and tasks, its scalability and generality are limited.
@@
Pattern-based relational extraction methods rely heavily on patterns written by linguistic experts, which are usually constructed from specific linguistic rules or regular expressions to capture fixed relational expressions in the text. This approach is very effective in domain-specific tasks, especially when the expected form of the relationship expression is relatively fixed. However, as the diversity of texts increases, the applicability of the fixed model begins to decline. For example, the sentences ‘France’s capital is Paris’ or ‘Paris, the capital of France’, which express the same relation, both require different models to be captured Sarawagi et al. (2008). To cope with this linguistic diversity, pattern-based approaches usually require the design of a large number of rules to cover the possible linguistic variants. These rules can be generated either by hand or by using some semi-automated tools, but either way, writing and maintaining them is very time-consuming and labour-intensive.
@@
Dependency Tree Based Relation Extraction method extracts more complex relations from sentences by analysing their dependency trees. The dependency tree is a commonly used syntactic structure representation in NLP, which represents the syntactic dependencies between words in a sentence. The advantage of the dependency tree approach is that it can handle complex grammatical structures, including long-distance dependencies and nested relationships. This ability makes the dependency tree-based approach outperform the simple pattern-based approach in processing complex sentences. Dependency tree-based rules usually involve multiple levels of conditions, e.g., the subject of a certain verb node should be a noun of a particular type, and the object of that verb needs to satisfy specific conditions. This fine-grained rule design allows the dependency tree-based approach to capture more complex types of relationships Sachan et al. (2020).
@@
  • Handling Complex Syntactic Structures: Dependency trees are able to capture complex relationships between words in a sentence, allowing the method to handle long-distance dependencies and nested structures, which is useful for relational extraction of complex sentences Tian et al. (2022).
@@
2.3.2 Deep Learning-based methods
Figure 4: Deep learning-based methods
With the rise of deep learning, the field of NLP has experienced a revolutionary change. Deep learning methods, shown in Figure 4, have greatly improved the performance of relation extraction by automatically learning to extract complex features from unlabelled data through neural network models. Compared with traditional rule-based approaches, deep learning is better able to adapt to the diversity of languages and performs particularly well when dealing with large-scale data Young et al. (2018).
@@ 2.3.2 Deep Learning-based methods
Convolutional Neural Network (CNN): The application of CNN in relationship extraction is mainly focused on processing sentence-level tasks. Using the sliding window, CNN can effectively extract local contextual information in a sentence, especially in capturing short-distance dependencies between words. Because CNN can process data in parallel, they are very suitable for training large-scale datasets and therefore have been widely adopted in practical applications Zeng et al. (2014).
Recurrent Neural Networks (RNNs) and Long Short-Term Memory Networks (LSTMs): RNNs and their variant LSTMs can capture temporal information in sequential data and are especially suitable for dealing with long-distance dependencies. LSTM solves the problem of gradient disappearance that occurs in traditional RNNs when dealing with long sequences through the introduction of memory units and the gating mechanism, which enables it to better remember the previous contextual information, thus improving the accuracy of relationship extraction Zhang and Wang (2015).
Attention-based models (e.g., Transformer): The introduction of the attention mechanism further improves the performance of deep learning models. The Transformer model is able to process each word while paying attention to the contextual information of all other words in the sentence through the self-attention mechanism, which makes the model perform much better in capturing complex relationships Vaswani (2017). Especially when dealing with long sentences and complex contexts, Transformer shows its advantages and has become one of the most popular deep learning models.
@@
Remote supervised learning is a weakly supervised approach to automatically generate annotated data by aligning text with an existing knowledge base. Remote/distance supervision assumes that if there is a known relationship between two entities in the knowledge base, then any sentence containing both entities may express that relationship. While this assumption simplifies the annotation process, it also introduces a large amount of noisy data, as not all sentences containing these entities express the relevant relationship Mintz et al. (2009); Yuan et al. (2019).
@@
3 Methodology
Figure 5: The pipeline of our LLM-Card method
In Figure 5, the names of large language models and the relationships between their applications or licences are extracted using LLM-Card methodology from academic literature using natural language processing techniques based on the dependency parsing approach. Dependency parsing is a technique that reveals syntactic dependencies between words in a sentence. It can systematically identify and extract key information describing large language models in the literature and convert it into structured relational data for further analysis.
3.1 Data Construction
3.1.1 Data collection and pre-processing
The first step of the study was to extract text from PDF documents from multiple sources and to ensure comprehensive coverage of the document’s content, including body text, footnotes, and graphical descriptions. The extracted text was subjected to disambiguation and lexical annotation using the spaCy library. Segmentation breaks down sentences into separate word units (tokens), while lexical annotation assigns each token a corresponding lexical tag (POS tag), such as noun (NOUN), verb (VERB), etc. As shown in Figure 10, in the sentence ‘BERT was released under the Apache licence 2.0’, where ‘BERT’ is tagged as a proper noun (PROPN), ‘was’ is tagged as an auxiliary verb (AUX), and ‘released’ is tagged as a verb (VERB). This process lays the foundation for dependency parsing.
3.1.2 Dependency Tree Construction

### v3

# AutoLLM-Card: Towards a Description and Landscape of Large Language Models

@@ Abstract.
With the rapid growth of the Natural Language Processing (NLP) field, a vast variety of Large Language Models (LLMs) continue to emerge for diverse NLP tasks. As more papers are published, researchers and developers face the challenge of information overload. Thus, developing a system that can automatically extract and organise key information about LLMs from academic papers is particularly important. The standard format for documenting information about LLMs is the LLM model card (LLM-Card). We propose a method for automatically generating LLM model cards from scientific publications. We use Named Entity Recognition (NER) and Relation Extraction (RE) methods that automatically extract key information about LLMs from the papers, helping researchers to access information about LLMs efficiently. These features include model licence, model name, and model application. With these features, we can form a model card for each paper. We processed 106 academic papers by defining three dictionaries – LLM’s name, licence, and application. 11,051 sentences were extracted through dictionary lookup, and the dataset was constructed through manual review of the final selection of 129 sentences with a link between the name and the licence, and 106 sentences with a link between the model name and the application. The resulting resource is relevant for LLM card illustrations using relational knowledge graphs. Our code and findings can contribute to automatic LLM card generation. Data and code in autoLLM-Card will be shared and freely available at https://github.com/shengwei-tian/dependency-parser-visualization
LLMs, Model Cards, NLP, Text Mining
copyright: acmcopyrightarticle: 4price: 15.00ccs: Computing methodologies Information extraction
1. Introduction
@@
As a major approach in natural language processing (NLP), language modelling has been widely studied for language understanding and generation in the past two decades, evolving from statistical language models to neural language models. Recently, Pre-trained Language Models (PLMs) have been proposed by pretraining Transformer models over large-scale corpora, showing strong capabilities in solving various NLP tasks.
@@
In recent years, major research institutes and technology companies have increased their investment in the development of LLMs, resulting in a series of groundbreaking models with the potential for a wide range of applications. OpenAI is a pioneer in this field. The GPT family (e.g., GPT-3 and GPT-4) has made significant advances in natural language generation, dialogue systems, and text comprehension, and is widely used in a variety of intelligent assistants, content creation tools, and automated services (Torfi et al., 2020). Meanwhile, the Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) and the subsequent T5 (Text-to-Text Transfer Transformer) (Raffel et al., 2020) models developed by Google research department have also made significant progress in tasks such as question-answer systems, text classification and information retrieval. The Roberta and BlenderBot series of models from Facebook (now Meta) have demonstrated powerful performance in social media data analysis and conversational AI (Liu et al., 2020).
@@
NLP technologies, driven by Large Language Models (LLMs), have skyrocketed and revolutionised the field of NLP, leading to breakthroughs in various tasks such as machine translation, text summarisation, and dialogue systems. However, these advances have been accompanied by the challenge of navigating the vast array of recent research findings. Each year, thousands of papers on LLMs are published, presenting new models, methods, and innovations. This leads to a phenomenon of information overload that makes it difficult for even the most experienced researchers to keep up with the latest developments.
Figure 1. autoLLM-card method demonstration
@@
We need a convenient and efficient tool that enables researchers and developers to quickly sift through and understand large numbers of models and their properties, i.e. the LLM model-card. Model cards often contain the metadata of the model providing some detailed technical documentation such as model parameters, usage, license, etc. (Singh et al., 2023; Liu et al., 2024)  [^1] https://huggingface.co/docs/hub/en/model-cards. Motivated by these needs, this work proposes a novel system that automatically extracts such relationships, as illustrated in Figure 1, the AutoLLM-Card. Our goal is to simplify the process of understanding key information about licence types, application areas, and more for researchers. By automating this process, researchers can gain a clearer understanding of the latest models, allowing them to save time to focus on innovation rather than tediously sifting through extensive literature.
@@
2. Related work
In this section, we talk about LLMs, Named Entity Recognition (NER), Relation Extraction (RE), and Model card generations.
2.1. Large Language Models
A pivotal moment in the history of LLMs came with the development of word embeddings, such as word2vec (Mikolov, 2013), which enabled models to learn distributed representations of words in a continuous vector space. This innovation laid the groundwork for more complex architectures like Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks, which were among the first to effectively model sequential data, capturing dependencies across different points in a text sequence.
Following BERT, the field of NLP has progressed rapidly with a series of more powerful models, notably OpenAI’s GPT family. GPT-2, released in 2019, significantly improves AI text generation and closes the gap with human writing. In 2020, GPT-3 set a new benchmark with 175 billion parameters for performing small or no samples for complex tasks such as code writing and dialogue generation (Brown, 2020). In 2022, ChatGPT based on GPT-3.5 further optimises dialogue performance for more natural interactions. GPT-4 in 2023 extends to multi-modal tasks, handling text and image inputs, showing potential for a wide range of applications in fields such as healthcare and law (OpenAI, 2023).
@@ 2.1. Large Language Models
2.2. Named Entity Recognition
@@
Named Entity Recognition (NER) is a subtask of information extraction in NLP, which is used to classify named entities into predefined categories such as names of people, organisations, places, etc. In the field of NLP, understanding these entities is crucial for many applications because they often contain the most important information in a text. NER is a bridge between unstructured text and structured data, enabling machines to sift through large amounts of textual information and extract valuable data in the form of classification. By precisely locating specific entities in a sea of words, NER changes the way we process and utilise textual data. In text mining, a named entity is a word or a phrase that clearly identifies an entity, an entity with similar properties, from a set of other items with similar properties. In the formulation of named entities, the word named limits the scope of entities that have one or more rigid designators that represent the referent. Typically, rigid designators often include proprietary names, but this depends on the domain of interest, which has reference words for objects as named entities (Sharnagat, 2014). For example, in the sentence: BERT was released under the Apache licence 2.0, the entity of licence is Apache licence and the entity of the model name is BERT.
Figure 2. The steps for named entities recognition
@@
As shown in Figure 2, NER can be broken down into several steps: Tokenisation before identifying entities, the text is split into tokens, which can be words, phrases, or even sentences. Entity identification involves recognising items such as capitalisation in licence (‘Apache 2.0’) or specific formats. Once entities are identified, they are grouped into predefined categories such as ‘model name’, ‘licence’ or ‘application’ in our task. This is typically achieved by machine learning models trained on annotated datasets. In the example “BERT was released under the Apache licence 2.0”, ‘BERT’ would be categorised as ‘model name’, while ‘Apache 2.0’ would be categorised as ‘licence’. The NER system usually considers the surrounding context to improve accuracy. After the initial identification and classification, post-processing may be performed to refine the results, which may involve resolving ambiguities, merging multi-tagged entities, or augmenting the entity data with a knowledge base (Jurafsky, 2000).
2.3. Relation Extraction
2.3.1. Rule-based methods
Figure 3. Rule-based methods for Relation Extractions
@@ 2.3.1. Rule-based methods
The rule-based relation extraction approach as shown in Figure 3 is one of the early techniques in the field of NLP, which relies on linguists and domain experts to manually write rules to extract relations from text based on an in-depth understanding of linguistic structures (Zhang et al., 2009). The core of this approach lies in the use of predefined linguistic patterns or syntactic structures to accurately match entities and their relationships in a text. While this approach excels in some specific domains and tasks, its scalability and generality are limited.
@@
Pattern-based relational extraction methods rely heavily on patterns written by linguistic experts, which are usually constructed from specific linguistic rules or regular expressions to capture fixed relational expressions in the text. This approach is very effective in domain-specific tasks, especially when the expected form of the relationship expression is relatively fixed. However, as the diversity of texts increases, the applicability of the fixed model begins to decline. For example, the sentences ‘France’s capital is Paris’ or ‘Paris, the capital of France’, which express the same relation, both require different models to be captured (Sarawagi et al., 2008). To cope with this linguistic diversity, pattern-based approaches usually require the design of a large number of rules to cover the possible linguistic variants. These rules can be generated either by hand or by using some semi-automated tools, but either way, writing and maintaining them is very time-consuming and labour-intensive.
@@
Dependency Tree Based Relation Extraction method extracts more complex relations from sentences by analysing their dependency trees. The dependency tree is a commonly used syntactic structure representation in NLP, which represents the syntactic dependencies between words in a sentence. The advantage of the dependency tree approach is that it can handle complex grammatical structures, including long-distance dependencies and nested relationships. This ability makes the dependency tree-based approach outperform the simple pattern-based approach in processing complex sentences. Dependency tree-based rules usually involve multiple levels of conditions, e.g., the subject of a certain verb node should be a noun of a particular type, and the object of that verb needs to satisfy specific conditions. This fine-grained rule design allows the dependency tree-based approach to capture more complex types of relationships (Sachan et al., 2020).
@@
  • Handling Complex Syntactic Structures: Dependency trees are able to capture complex relationships between words in a sentence, allowing the method to handle long-distance dependencies and nested structures, which is useful for relational extraction of complex sentences (Tian et al., 2022).
@@
2.3.2. Deep Learning-based methods
Figure 4. Deep learning-based methods
With the rise of deep learning (DL), the field of NLP has experienced a revolutionary change. Deep learning methods, shown in Figure 4, have greatly improved the performance of relation extraction by automatically learning to extract complex features from unlabelled data through neural network models. Compared with traditional rule-based approaches, deep learning is better able to adapt to the diversity of languages and performs particularly well when dealing with large-scale data (Young et al., 2018).
@@ 2.3.2. Deep Learning-based methods
Convolutional Neural Network (CNN): The application of CNN in relationship extraction is mainly focused on processing sentence-level tasks. Using the sliding window, CNN can effectively extract local contextual information in a sentence, especially in capturing short-distance dependencies between words. Because CNN can process data in parallel, they are very suitable for training large-scale datasets and therefore have been widely adopted in practical applications (Zeng et al., 2014).
Recurrent Neural Networks (RNNs) and Long Short-Term Memory Networks (LSTMs): RNNs and their variant LSTMs can capture temporal information in sequential data and are especially suitable for dealing with long-distance dependencies. LSTM solves the problem of gradient disappearance that occurs in traditional RNNs when dealing with long sequences through the introduction of memory units and the gating mechanism, which enables it to better remember the previous contextual information, thus improving the accuracy of relationship extraction (Zhang and Wang, 2015).
Attention-based models (e.g., Transformer): The introduction of the attention mechanism further improves the performance of deep learning models. The Transformer model is able to process each word while paying attention to the contextual information of all other words in the sentence through the self-attention mechanism, which makes the model perform much better in capturing complex relationships (Vaswani, 2017). Especially when dealing with long sentences and complex contexts, Transformer shows its advantages and has become one of the most popular deep learning models.
@@
Remote supervised learning is a weakly supervised approach to automatically generate annotated data by aligning text with an existing knowledge base. Remote/distance supervision assumes that if there is a known relationship between two entities in the knowledge base, any sentence containing both entities may express that relationship. While this assumption simplifies the annotation process, it also introduces a large amount of noisy data, as not all sentences containing these entities express the relevant relationship (Mintz et al., 2009; Yuan et al., 2019).
@@
2.4. Model Card Generation
In this subsection, we introduce the very few work available, to our best knowledge, on LLM model card generation.
The first work is from (Singh et al., 2023) who introduced a dataset to facilitate model training on model card generation. The data set contains 500 QA pairs for 25 learning models. The metadata aims to cover the model model training configurations, datasets used, biases, architecture details. Their investigation on ChatGPT, LLaMa, and Galactica identified the performance gap between these models and humans on factual text responses.
The other two very recent works from this year are from (Liu et al., 2024) and (Yang et al., 2024). (Liu et al., 2024) further extended the dataset size to cover 4.8k model cards and 1.4k data cards using their two-step retrieval processing (CardBench and CardGen). Instead, the work from (Yang et al., 2024) (Report Cards) features on three criteria they designed: specificity, faithfulness, and interpretability. In addition, they proposed a baseline algorithm that does not need human supervision.
In comparison to these works, we propose to use different methodology by leveraging dependency parsing to extract model names and their (applications, licences).
3. AutoLLM-Card Methodology
Figure 5. The pipeline of our AutoLLM-Card method
In Figure 5, the names of large language models and the relationships between their applications or licences are extracted using AutoLLM-Card methodology from academic literature using natural language processing techniques based on the dependency parsing approach. Dependency parsing is a technique that reveals syntactic dependencies between words in a sentence. It can systematically identify and extract key information describing large language models in the literature and convert it into structured relational data for further analysis.
Figure 6. Dependency tree for the active voice processing sentence: GPT-4 enhances text generation with more accurate and context-aware outputs.
Figure 7. Dependency Tree Based Rules on Passive and Active Voice Syntax.
Figure 8. Pseudo Code for Algorithm Illustration
3.1. Data Construction and Relation Extraction
3.1.1. Data collection and pre-processing
The first step of the study was to extract text from PDF documents from multiple sources and to ensure comprehensive coverage of the document’s content, including body text, footnotes, and graphical descriptions. We removed some unnecessary structural information from the PDF document. The extracted text was subjected to disambiguation and lexical annotation using the spaCy library (Honnibal et al., 2020)  [^2] https://spacy.io/. Segmentation breaks down sentences into separate word units (tokens), while lexical annotation assigns each token a corresponding lexical tag (POS tag), such as noun (NOUN), verb (VERB), preposition (ADP), adverb (ADV), etc. SpaCy will use these POS tags to extract dependency relation. As shown in Figure 12, in the sentence ‘BERT was released under the Apache licence 2.0’, where ‘BERT’ is tagged as a proper noun (PROPN), ‘was’ is tagged as an auxiliary verb (AUX), and ‘released’ is tagged as a verb (VERB). We removed all words “the” because we do not think they are useful for the relation extraction task. Next, we introduce the relation extraction using dependency trees.
3.1.2. Dependency Tree Construction using spaCy